OpenStudy (anonymous):

Because int(1/x^2, x = -2 .. -1) the result is 1/2?

OpenStudy (anonymous):

\[\int_{-2}^{-1}\frac{dx}{x^2}\]??

OpenStudy (anonymous):

yeah, what i wrote

OpenStudy (anonymous):

anti derivative is \(-\frac{1}{x}\)

OpenStudy (anonymous):

plug in \(-2\) then plug in \(-1\) and subtract

OpenStudy (anonymous):

I did not

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[-\frac{1}{x}\bigg|_{-2}^{-1}\] is the usual notation but it means \[ f(-2)-F(-1)\]

OpenStudy (anonymous):

you get \[-\frac{1}{-2}-(-\frac{1}{-1})\] sort out the minus signs and you will get it

OpenStudy (anonymous):

oops wrong, i had that backwards!

OpenStudy (anonymous):

\[-\frac{1}{-1}-(-\frac{1}{-2})\]

OpenStudy (anonymous):

so yeah, \(1-\frac{1}{2}=\frac{1}{2}\) is the answer

OpenStudy (anonymous):

Now I could understand my friend, thank you ... you use facebook?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

whatsapp, skype?